OpenStudy (anonymous):

how to solve this - ( m + 2 ) = 3 + m ?

myininaya (myininaya):

Do you know the distributive property?

myininaya (myininaya):

a(b+c)=ab+ac

myininaya (myininaya):

You should apply that on your left hand side of the equation.

OpenStudy (anonymous):

yes kind of

myininaya (myininaya):

So can you tell me what you get after using the distributive property on: -(m+2)?

OpenStudy (anonymous):

hold it

OpenStudy (anonymous):

so the M is one right?

myininaya (myininaya):

No. Let's go in steps. I'm trying to get you distribute right now -1(m+2)=?

OpenStudy (anonymous):

1(m+2)

OpenStudy (anonymous):

i kind of forgot everything already O;

myininaya (myininaya):